5x^(2)+55x+150= an u show me the solution of this problem and the work and the answer to this question

Answer by checkley79(3341)

5x^(2)+55x+150=0
5(X^2+11X+30)=0
5(X+5)(X+6)=0
X+5=0
X=-5 ANS.
X+6=0
X=-6 ANS.